### v1

# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

Abstract.
Voice Authentication Systems (VAS) use unique vocal characteristics for verification. They are increasingly integrated into high-security sectors such as banking and healthcare. Despite their improvements using deep learning, they face severe vulnerabilities from sophisticated threats like deepfakes and adversarial attacks. The emergence of realistic voice cloning complicates detection, as systems struggle to distinguish authentic from synthetic audio. While anti-spoofing countermeasures (CMs) exist to mitigate these risks, many rely on static detection models that can be bypassed by novel adversarial methods, leaving a critical security gap. To demonstrate this vulnerability, we propose the Spectral Masking and Interpolation Attack (SMIA), a novel method that strategically manipulates inaudible frequency regions of AI-generated audio. By altering the voice in imperceptible zones to the human ear, SMIA creates adversarial samples that sound authentic while deceiving CMs. We conducted a comprehensive evaluation of our attack against state-of-the-art (SOTA) models across multiple tasks, under simulated real-world conditions. SMIA achieved a strong attack success rate (ASR) of at least 82% against combined VAS/CM systems, at least 97.5% against standalone speaker verification systems, and 100% against countermeasures. These findings conclusively demonstrate that current security postures are insufficient against adaptive adversarial attacks. This work highlights the urgent need for a paradigm shift toward next-generation defenses that employ dynamic, context-aware frameworks capable of evolving with the threat landscape.
Machine Learning Security , Voice Authentication, Anti-Spoofing, Voice Cloning, Adversarial Spoofing Attacks
$c$s: Security and privacy Biometricsccs: Computing methodologies Artificial intelligenceccs: Security and privacy Systems security
1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (Singh et al., 2012; Algabri et al., 2017; Rose, 2006). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
The global voice biometrics market was approximately USD 2.5 billion in 2023 and is projected to reach between USD 11.5 billion and USD 15.7 billion by 2032, growing at a compound annual growth rate (CAGR) ranging from 18.7% to 27.5% during this period (Allied Market Research, 2023). This growth is driven by the increased demand for robust fraud detection and prevention systems, particularly within the financial services sector (Markets and Markets, 2023; Straits Research, 2023; Fortune Business Insights, 2024).
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (Renz et al., 2023; Springer, 2023; Techlocity, 2021; Khan and Zahid, 2010).
Recent advancements in deep learning have revolutionized voice authentication systems (VASs), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (Lan et al., 2022).
In parallel, deep learning has also transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (Liao et al., 2024; Qin et al., 2023; Amazon Web Services, 2025) and voice conversion (VC) (Li et al., 2023; Huang et al., 2024) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass Voice Authentication Systems (VAS) (Shonesy, 2015; Alali and Theodorakopoulos, 2025).
To combat these deepfake attacks, various anti-spoofing countermeasures (CMs), like RawNet2 (Tak et al., 2021b) or RawGAT-ST (Tak et al., 2021a), have been developed to distinguish between genuine and synthetic speech by identifying subtle artifacts that differentiate the two. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: to bypass the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (Hai et al., 2023; Kassis and Hengartner, 2023).
A compelling example is SiFDetectCracker (Hai et al., 2023), an adversarial attack that can deceive state-of-the-art CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, the effectiveness of these attacks is significantly reduced in black-box transferability tests, and their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are not only effective but also transferable across different systems in real-world scenarios.
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. A pivotal work by Kassis et al. (Kassis and Hengartner, 2023) demonstrated the feasibility of a practical, model-agnostic attack that bypasses these CMs not through query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (Amazon, 2025) and RawDarts (Ge et al., 2021). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of state-of-the-art, hardened VA systems. Moreover, designated-app attack vectors often require privileged device modifications, such as rooting a phone to inject audio directly, which poses a substantial barrier for real-world exploitation in security-critical contexts.
Our main contributions of this paper are summarized as follows:
  • We introduce the Spectral Masking and Interpolation Attack (SMIA). This novel black-box adversarial spoofing attack exploits inaudible regions of AI-generated audio to bypass both VAS and CMs.
  • We conduct extensive evaluations of SMIA against widely adopted open-source VAS (Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018)), commercial platforms (Microsoft Azure SV (Microsoft, [n. d.])), and state-of-the-art CMs (RawNet2 (Tak et al., 2021b), RawGAT-ST (Tak et al., 2021a), and RawPC-Darts (Ge et al., 2021)).
  • Our results demonstrate that SMIA significantly outperforms current state-of-the-art attacks by addressing critical methodological gaps, thereby creating a more robust and effective threat.
Unlike prior work, which explored TTS-based attacks, targeted voice authentication systems or countermeasures in isolation, or relied on transferability-based attacks with limited effectiveness, we present a black-box attack that simultaneously compromises both. Our evaluations span a wide range of models, tasks, and real-world scenarios, demonstrating the consistency and severity of this vulnerability. The results reveal that VAS are significantly more fragile than previously assumed, underscoring the urgent need for stronger defenses and providing a solid foundation for future research in biometric security.
This paper is organized as follows. Section 2 provides an overview of voice authentication systems and their associated threats, related works, and the threat model. Section 3 outlines our methodology and describes the SMIA workflow. Section 4 details the experimental setup, including the datasets, models, and configurations used. Section 5 presents the evaluation of our attack. Section 6 reports the results of our experiments, highlighting the effectiveness and limitations of the studied approaches. Section 7 provides an ablation study, analyzing the impact of different components and design choices. Section 8 discusses the implications of our findings and covers the challenges and open questions in securing voice authentication systems. Section 9 outlines future work, suggesting directions to improve system robustness and defense mechanisms. Finally, Section 10 concludes the paper by summarizing the main insights and contributions.
2. Background and Related Works
2.1. List of Acronyms
For clarity, Table 1 defines the acronyms used throughout this document.
Table 1. List of Acronyms
Acronym Definition
ABTN Adversarial Biometrics Transformation Network
ASR Attack Success Rate
CAGR Compound Annual Growth Rate
CM Countermeasure
CQCCs Constant-Q Cepstral Coefficients
CSI Closed-Set Identification
DNN Deep Neural Network
EER Equal Error Rate
FGSM Fast Gradient Sign Method
GAN Generative Adversarial Network
GAT Graph Attention Network
GMMs Gaussian Mixture Models
LA Logical Access
LCNN Light Convolutional Neural Network
MFCCs Mel-Frequency Cepstral Coefficients
NAS Neural Architecture Search
OSI Open-Set Identification
PGD Projected Gradient Descent
ResNet Residual Networks
SiFs Speaker-irrelevant Features
SMIA Spectral Masking and Interpolation Attack
SOTA State-of-the-Art
SRS Speaker Recognition Systems
SSL Self-Supervised Learning
STFT Short-Time Fourier Transform
SV Speaker Verification
SVMs Support Vector Machines
TDNN Time Delay Neural Network
TTS Text-to-Speech
VA Voice Authentication
VAS Voice Authentication System
VC Voice Conversion
2.2. Speaker Recognition Systems (SRS)
SRS identifies individuals through their vocal features. The authentication process typically unfolds in three key phases, as shown in Figure. 1.
  1. (1)
    First, during the training phase, the system learns voice characteristics from a large corpus to develop robust voice authentication models.
  2. (2)
    Next, in the enrollment phase, a user’s voice is recorded as they speak specific phrases, and the extracted embeddings are stored.
  3. (3)
    Finally, during the recognition phase, the system compares embeddings from the newly captured voice sample to the enrolled embeddings to confirm identity.
Figure 1. Speaker Recognition System Overview. 1) Training phase. 2) Enrollment phase. 3) Verification phase.
2.3. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (Patel and Patil, 2015; De Leon et al., 2017). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has since shifted towards deep learning methods, particularly convolutional and recurrent architectures designed to learn discriminative spectral and temporal patterns directly from audio spectrograms (Lavrentyeva et al., 2019). More recently, the state-of-the-art has been driven by Self-Supervised Learning (SSL) models used as powerful front-end feature extractors. Models such as Wav2Vec2.0 (Baevski et al., 2020) and especially WavLM (Chen et al., 2022) have become dominant, with WavLM showing superior performance due to its denoising and masked speech prediction pre-training objectives, which capture fine-grained speaker-related artifacts commonly found in deepfakes. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (Lavrentyeva et al., 2017) or the AASIST model (Jung et al., 2022), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (Tak et al., 2021b) demonstrated strong performance in the ASVspoof 2019 challenge (et al., 2019), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as state-of-the-art alternatives. Raw-PC-DARTS (Ge et al., 2021) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, Raw-GAT-ST (Tak et al., 2021a) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
2.4. Attacks Taxonomies
Attacks can be categorized based on the attacker’s goal to:
  1. (1)
    Targeted Attacks: Forcing the system to recognize the input as a specific enrolled speaker.
  2. (2)
    Untargeted Attacks: Causing misclassification as any incorrect speaker without control over the output identity.
These attacks can also be categorized based on the attacker’s knowledge:
  • White-box Attack: The attacker fully knows the system, including its architecture, parameters, training data, and feature extraction methods. This enables precise gradient-based crafting of adversarial examples.
  • Grey-box Attack: The attacker possesses partial knowledge, such as the feature representation or model architecture, but lacks access to full internals like trained weights. Attacks typically exploit transferability or inferred system behavior.
  • Black-box Attack: The attacker lacks any internal knowledge and treats the system like an oracle. They rely on input-output queries. Attack methods usually involve optimizing queries or using transfer attacks with surrogate models.
2.5. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (Ren et al., 2019, 2020; Kim et al., 2020; Kaneko and Kameoka, 2018; Kameoka et al., 2018; Qian et al., 2019; Li et al., 2023; Qin et al., 2023; Wang et al., 2024). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (Almutairi and Elgibreen, 2022).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (Shonesy, 2015). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: state-of-the-art speaker recognition systems failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (Alali and Theodorakopoulos, 2025) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and speaker recognition systems remain vulnerable when exposed to partial speech generated through zero-shot text-to-speech and voice conversion models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
2.6. Adversarial Spoofing Attacks
These attacks typically involve adding subtle perturbations to AI-generated audio, crafted by speech synthesis models, to fool CMs or both the VAS and CMs.
2.6.1. White-box Attacks
Early work by Liu et al. (Liu et al., 2019) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to countermeasure (CM) systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51
In a related study, Alanis et al. (Gomez-Alanis et al., 2022b) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.
2.6.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. proposed SiFDetectCracker, a framework targeting fake voice detection systems (CMs). Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), RawPC-Darts (Ge et al., 2021), and Deep4SNet (Ballesteros et al., 2021), using the ASVspoof 2019 dataset (et al., 2019). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (Kassis and Hengartner, 2023) presents a novel methodology for bypassing voice authentication (VA) systems, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (Ge et al., 2021) and RawGAT-ST (Tak et al., 2021a), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
Table 2. Summary of Key Deepfake and Adversarial Spoofing Attacks on Speaker Recognition Systems
Paper Year Tasks Attack Type System Target
Deepfake Voice Attacks
Saxena et al. (Shonesy, 2015) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (Alali and Theodorakopoulos, 2025) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
Adversarial Spoofing Attacks
Liu et al. (Liu et al., 2019) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (Gomez-Alanis et al., 2022b) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (Hai et al., 2023) 2023 CM Black-box CM
Kassis et al. (Kassis and Hengartner, 2023) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
2.7. Threat Model
Our threat model considers an attacker whose primary goal is to gain unauthorized access to a victim’s account on a Speaker Verification (SV) system by mimicking their voice. The success of the attack is defined by a single successful authentication event.
The attacker operates under a black-box assumption, meaning they have no internal knowledge of the system’s architecture, including the specific Voice Authentication System (VAS) or Countermeasure (CM) models employed. The attacker’s capabilities are as follows:
  • They can collect voice samples of the victim from publicly available sources, requiring a minimum of 10 seconds of audio to initiate the attack.
  • They can generate a spoofed audio sample using a voice-cloning tool (e.g., Fish Speech APIs (Liao et al., 2024), methods used in ASVspoof 2019 (et al., 2019), or any similar tools).
  • The attacker can repeatedly submit the generated audio to the system’s authentication endpoint and observe the binary pass/fail outcome. This feedback is used to iteratively optimize the malicious audio sample to bypass both the VAS and CM defenses.
The attacker’s actions are considered stealthy, as a successful login does not interfere with or alert the legitimate user. The attack is applicable to both text-dependent and text-independent SV systems, though text-dependent systems present a higher barrier, requiring the attacker to also know the specific pass-phrase for the authentication session.
3. Methodology
A combined voice authentication and anti-spoofing system processes a given voice sample $x\in\mathbb{R}^{m}$ to perform two key tasks: speaker verification (SV) and spoofing detection (CM). The system grants access only when the sample is verified as belonging to the claimed identity and determined to be a genuine, live utterance.
This integrated system can be formally defined with a speaker verification model $g(x,e)$ and an anti-spoofing model $f(x)$. For a given input $x$ and an enrolled embeddings $e$, the system’s decision $S(x,e)$ is given by:
(1) $S(x,e)=\begin{cases}\text{accept}&\text{if }g(x,e)=\text{accept}\land f(x)=\text{real}\\ \text{reject}&\text{otherwise}\end{cases}$
The objective of our attack is to generate an adversarial sample $x^{\prime}$ from an initial spoofed utterance $x$ that bypasses this combined system. This requires crafting an $x^{\prime}$ that is both accepted by the speaker verification model and misclassified as real by the anti-spoofing model. In short, our goal is to ensure $S(x^{\prime},e)=\text{accept}$.
In this section, we first present Spectral Masking and Interpolation Attack (SMIA), which manipulates these features to bypass the countermeasure while preserving the speaker characteristics needed to fool the verification model, before delving into its details.
3.1. Overview
We propose a novel black-box adversarial attack framework designed to circumvent both VAS and their corresponding CMs. The attack operates by combining high-fidelity voice synthesis with an adaptive, iterative perturbation process.
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—you can employ Fish Speech APIs (Liao et al., 2024), ASVSpoof challenge tools (et al., 2019), or any modern voice cloning system.
Figure 2. Our Spectral Masking & Interpolation Attack (SMIA) that targets both Voice Authentication and Anti-Spoofing Systems.
Second, this synthesized audio serves as the input to our core attack mechanism: the Spectral Masking and Iterative Attack (SMIA) system. SMIA operates within an iterative loop where it strategically modifies the audio using our proposed Spectral Masking and Interpolation technique. The resulting adversarial audio is submitted to the target system (VAS or CM), and the prediction outcome is used as feedback to update the SMIA parameters. This process is repeated, progressively refining the audio until it successfully deceives the target model.
The core of the SMIA technique is a stochastic perturbation method. In each iteration, the attack applies a probabilistic masking operation to low-energy (i.e., quiet) regions of the audio’s frequency spectrum. This strategy ensures the perturbations are perceptually inconspicuous, making the attack virtually undetectable by human listeners. Furthermore, because these modifications are subtle and distributed across the spectrum rather than being concentrated as predictable noise, they are inherently resilient to conventional audio filtering or preprocessing defenses. While each perturbation is minor, they accumulate over iterations to create a significant distortion in the feature space recognized by the models, leading to a successful attack.
Our attack framework is thus composed of two integral components:
  • Iterative Black-Box Optimization: A feedback-driven loop that queries the target model and uses its decision outcome to adaptively update the params to guide the generation of the adversarial audio.
  • Spectral Masking and Interpolation: A stealthy perturbation method that introduces distortions by targeting perceptually insignificant spectral regions, thereby preserving the audio’s naturalness.
Together, these components create a robust and adaptive feedback-driven process that dynamically refines its adversarial strategy, making the attack highly effective and difficult to defend against.
3.2. Iterative Black-Box Optimization
Our attack is operationalized as a black-box iterative search designed to find an adversarial audio sample that simultaneously bypasses both the voice authentication (VAS) and anti-spoofing (CM) systems. The core of this process, detailed in Algorithm 1, employs a multi-modal perturbation strategy. Rather than relying on a single modification technique, the algorithm systematically cycles through a set of distinct perturbation modes: Interpolate, Masking, and a Hybrid approach. For each mode, it iteratively applies subtle modifications to a clone of the original spoofed audio. Based on the feedback from each attempt, the loop updates its core parameters ($T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}$) to guide the subsequent search. The goal at each step is to produce a variant that the VAS accepts as the target speaker and the CM classifies as bona fide (i.e., not spoofed). This adaptive, multi-pronged search increases the likelihood of discovering a successful adversarial example without requiring any internal knowledge of the target models.
Algorithm 1 Multi-Modal Adaptive Adversarial Attack
1:Enrollment embedding $E$, initial attacker audio $x_{orig}$
2:Max iterations per mode NN
[…]
4:for each modemode in ModesModes do
5:  $x\leftarrow x_{orig}$
6:  Init perturbation params $P\leftarrow\{T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}\}$
7:  for $i\leftarrow 0$ to N1N-1 do
8:   $x_{adv}\leftarrow\text{ApplyPerturbation}(x,mode,P)$
9:   $vas\_label\leftarrow M_{VAS}(x_{adv},E)$
10:   if $vas\_label=\text{`Accept'}$ then
11:     $cms\_label\leftarrow M_{CMS}(x_{adv})$
12:     if $cms\_label=\text{`Bona Fide'}$ then
13:      return $(x_{adv},\textbf{Success})$
14:     else
15:      $P\leftarrow\text{IncreasePerturbation}(P)$
16:     end if
17:   else
18:     $P\leftarrow\text{DecreasePerturbation}(P)$
19:   end if
20:  end for
21:end for
22:return $(x_{orig},\textbf{Failure: Max iterations reached for all modes})$
3.3. Spectral Masking and Interpolation
At the core of our attack is a sophisticated spectral perturbation module, ApplyPerturbation. The primary function of this module is to introduce subtle, controlled distortions into an audio signal’s spectrogram. By design, it operates exclusively on low-energy (quiet) time-frequency bins, which are perceptually less significant to the human ear. This strategic targeting is controlled by five key parameters (as shown in Algorithm 2):
  • $T\_{\text{db}}$ A decibel threshold used to identify the quiet parts of the spectrogram.
  • $\gamma$ and $\sigma$ Parameters that control the random selection of bins to be perturbed within the quiet regions.
  • n_fft and hop_length STFT (Short-Time Fourier Transform) settings that define the resolution of the spectrogram grid, which directly influences the interpolation process.
This approach ensures that the resulting adversarial audio remains natural-sounding and free of obvious artifacts. The module employs three primary techniques to manipulate the audio’s underlying acoustic features: masking, interpolation, and a hybrid approach.
Algorithm 2 Spectral Perturbation Module
1:Audio signal $x$, modemode
2:Perturbation params $P=\{T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}\}$
[…]
9:if $mode=\text{`interpolate'}$ or $mode=\text{`hybrid'}$ then
10:  $\text{mag\_out}\leftarrow\text{Interpolate}(\text{mag\_out},\text{final\_mask})$
11:end if
12:if $mode=\text{`mask'}$ or $mode=\text{`hybrid'}$ then
13:  $\text{mag\_out}[\text{final\_mask}]\leftarrow 0$
14:end if
[…]
16:return $x_{adv}$
3.3.1. Perturbation Modes
Masking
In this mode, the module takes a straightforward approach: it resets the magnitude of the selected quiet and randomized time-frequency bins to zero. This effectively silences these specific parts of the signal, creating a subtle distortion by removing minimal acoustic information.
Interpolation
This mode aims to replace the targeted quiet bins with new values that are consistent with the surrounding stable, non-quiet parts of the signal. For each frequency channel in the spectrogram, the module performs the following steps:
  1. (1)
    It identifies all the reliable, high-energy (non-quiet) points over time, using them as anchors.
  2. (2)
    It constructs a linear interpolation function based on these anchor points.
  3. (3)
    It then uses this function to predict what the values should have been in the targeted quiet locations.
Essentially, it reconstructs the quiet segments based on the audio’s natural temporal contour within that specific frequency. The value $y$ for a target point $x$ between two known anchor points $(x_{0},y_{0})$ and $(x_{1},y_{1})$ is calculated using the standard linear interpolation formula:
(2) $y(x)=y_{0}+(x-x_{0})\frac{y_{1}-y_{0}}{x_{1}-x_{0}}$
This process ensures that the filled-in values are spectrally smooth and plausible, making the alteration difficult to perceive.
Hybrid
The hybrid mode offers the most flexibility by combining both masking and interpolation. It applies a mix of both techniques to the targeted bins, creating a more complex and varied spectral perturbation. This allows the attack to leverage the distinct advantages of each method simultaneously.
3.3.2. Efficacy Against VAS and CMS
This dual-operation approach is highly effective because it simultaneously targets the distinct vulnerabilities of both VAS and CMs. The primary challenge is to bypass the CMS without sacrificing the convincing vocal identity needed to fool the VAS.
  • Deceiving the Voice Authentication System (VAS): The initial spoofed audio, generated by Text-to-Speech (TTS) or Voice Conversion (VC) tools, is already designed to mimic the target speaker’s identity. Our main goal is to preserve this biometric similarity during the attack. By confining perturbations to the low-energy (quiet) regions, we alter the signal’s subtle characteristics, which are scrutinized by a CMS, while leaving the dominant, identity-defining acoustic patterns largely intact. This surgical approach ensures the audio retains the core vocal signature required for the VAS to accept it as the legitimate user.
  • Evading the Countermeasure System (CM): A CM is trained to detect artificial signals by identifying unnatural artifacts, such as vocoder noise or phase discontinuities. Our interpolation technique directly counters this. By filling quiet regions with spectrally consistent data derived from the surrounding audio, we generate a spectrogram that appears smooth and natural. This process removes or masks many of the sharp, anomalous features that a CMS typically flags as evidence of a spoof, allowing the audio to be classified as bona fide.
3.4. Design Justification
  • Quiet Region Targeting: Perturbations in quiet regions are less perceptible to human listeners and allow for stealthy adversarial modifications.
  • Adaptive Strategy: Dynamic adjustment of perturbation parameters improves attack success while maintaining signal intelligibility. Our approach explicitly tests both anti-spoofing and voice authentication systems (VAS). When perturbations are overly strong, similarity with the VAS is lost; when too minimal, the anti-spoofing system cannot be bypassed. This adaptive strategy is well-suited to balance these conflicting objectives.
  • Randomized Selection: Introducing a random element to the selection of bins offers two key advantages. First, it provides greater flexibility regarding the regions of the spectrogram that are modified. Second, by constantly changing the perturbation pattern, it makes the attack significantly harder to detect, as there is no fixed, predictable signature for a defense system to learn.
4. Experimental Setup
Our experimental setup leverages the LibriSpeech (Panayotov et al., 2015) and ASVSpoof2019 (et al., 2019) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), and RawPC-Darts(Ge et al., 2021). For the voice authentication stage, we employ three distinct systems: X-Vectors (Snyder et al., 2018), DeepSpeaker (Li et al., 2017), and the commercial Azure Speaker Verification API (Microsoft, [n. d.]). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations:
  • Combined systems integrating both anti-spoofing and voice authentication.
  • Anti-spoofing systems in isolation.
  • Voice Authentication Systems in isolation.
4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (et al., 2019) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (Panayotov et al., 2015) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
4.1.1. ASVSpoof 2019
The ASVSpoof 2019 dataset is the primary benchmark for evaluating spoofing countermeasures. We use the Logical Access (LA) portion, whose partitions are statistically summarized in Table 3. To create a robust and challenging test set for our experiments against anti-spoofing systems, we curated a specific subset from the official Evaluation partition. As shown in the table, we selected 100 voice samples for each of the 13 advanced spoofing attack types (A07–A19), resulting in a balanced test set of 1,300 spoofed samples for our analysis.
Table 3. Statistics of the ASVSpoof 2019 LA Dataset and Experimental Subset.
Partition Speakers (M/F) Total Speakers Total Utt. Selected (per Attack) Total Selected
Training 8 / 12 20 25,380
Development 8 / 12 20 24,844
Evaluation 29 / 39 68 71,237 100 1,300
4.1.2. LibriSpeech
For our end-to-end experiments evaluating the full pipeline, we used the dev-clean and test-clean subsets of the LibriSpeech ASR corpus. As detailed in Table 4, these subsets provide high-quality audio from a total of 80 unique speakers, evenly distributed between males and females. To simulate a realistic user verification scenario, we selected a specific number of samples for each speaker. From each of the 80 speakers, we designated 5 utterances for enrollment to create their voiceprint and a separate set of 5 utterances for verification trials. This resulted in a total of 400 selected audio files from each subset.
Table 4. Statistics of the LibriSpeech Subsets and Experimental Selection.
Subset Speakers (M/F) Total Speakers Utterances Selected (per User) Total Selected
dev-clean 20 / 20 40 2,703 10 (5 enroll, 5 verify) 400
test-clean 20 / 20 40 2,620 10 (5 enroll, 5 verify) 400
4.2. Authentication Models
To ensure a comprehensive evaluation, we test our attack against a diverse set of three state-of-the-art (SOTA) Speaker Verification (SV) systems. This set includes two well-known open-source models and one commercial black-box API, allowing us to assess performance on both accessible academic benchmarks and proprietary, real-world systems.
  • X-Vectors (Snyder et al., 2018): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (Ravanelli et al., 2021) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (Li et al., 2017)  [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (Microsoft, [n. d.])  [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (et al., 2019). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (Tak et al., 2021b)  [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (et al., 2019). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (Tak et al., 2021a)  [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (et al., 2019).
  • RawPC-DARTS (Ge et al., 2021)  [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (et al., 2019), making it a critical system to evaluate against.
5. Evaluation Methodology
To comprehensively assess the effectiveness of our proposed attack, we designed a series of rigorous experiments targeting SOTA VAS and CMs. Our methodology is structured to benchmark performance against prior work, evaluate end-to-end systems, and test robustness in realistic scenarios.
5.1. End-to-End System Evaluation on ASVspoof 2019
To ensure our methodology is grounded in a standard framework, we conduct our experiments on the ASVspoof 2019 dataset (et al., 2019). This dataset is a well-established and widely recognized benchmark for research on voice spoofing and countermeasures. While we do not perform a direct comparison with existing literature in this study, our use of this dataset provides a crucial baseline. This approach allows future work to easily and effectively utilize our results for a direct comparison with our attack. In this setup, we use the genuine bona fide audio samples for speaker enrollment and the dataset’s pre-existing spoofed audio files as the base for our perturbations.
5.2. End-to-End System Evaluation on LibriSpeech
Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (Panayotov et al., 2015). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (Liao et al., 2024). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
5.3. Targeted System Evaluations
To gain deeper insights into the vulnerabilities of individual system components, we conduct targeted experiments against CMs and VAS separately.
  • Attacks on Anti-Spoofing Systems using LibriSpeech: We first perform a dedicated experiment targeting only the CMs. As described above, these test audios were generated from a pair combination of different prompts and LibriSpeech audios (Panayotov et al., 2015) that passed to Fish Speech APIs (Liao et al., 2024), then passed to our SMIA attack to generate adversarial audios that tested against different CMs to test their vulnerability.
  • Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019: To ensure a direct and fair comparison with prior work, we conduct a second experiment targeting CMs using the ASVspoof 2019 dataset (et al., 2019). We used the spoofed audios from the ASVspoof 2019 dataset (et al., 2019), then passed them to our SMIA framework to test them against different CMS. This methodology allows for a clear performance benchmark against the results reported by both Kassis et al. (Kassis and Hengartner, 2023) and Hai et al. (Hai et al., 2023).
  • Attacks on Voice Authentication Systems: We conduct an extensive evaluation to demonstrate the significant vulnerability of VAS. Using the LibriSpeech (Panayotov et al., 2015) and Fish Speech (Liao et al., 2024), we test our generated audio against the VAS models across three distinct biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
  • Computational Efficiency Evaluation: To provide a complete picture, we evaluated SMIA’s computational efficiency against SiFDetectCracker (Hai et al., 2023). We measured both the number of iterations and the total time needed to carry out a successful attack. This analysis confirms our framework’s practical feasibility, highlighting its significant speed advantage over prior work.
  • Robustness in Real-World Scenarios: An effective attack must be robust to the distortions introduced by real-world transmission channels. To evaluate our method’s practicality, we test the generated audio against the VAS systems in two simulated environments: over-the-air, which mimics audio being played through a speaker and recorded by a microphone, and over-the-line, which simulates the effects of a VoIP codec. This demonstrates the attack’s applicability beyond controlled digital environments.
5.4. Stealth and Detectability Analysis
To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative analysis. We show that while adversarial audio generated by methods, like SiFDetectCracker (Hai et al., 2023) can be easily identified and neutralized by a simple high-pass filter, our attack audio remains effective. This highlights our method’s superior stealth and its ability to bypass rudimentary digital forensic defenses.
5.5. Ablation Study
To understand which parts of our attack contribute most to its success, we conducted two specific ablation studies.
  1. (1)
    Attack Modes: The first study focuses on our perturbation strategies. We systematically compared the performance of the three distinct modes—Masking, Interpolation, and Hybrid—to determine which method is most effective at bypassing the target systems.
  2. (2)
    Randomness: The second study investigates the importance of our stochastic approach. We evaluated the attack’s performance both with and without the random selection of quiet bins. This analysis clarifies how randomness contributes to the attack’s flexibility and ability to evade detection.
  3. (3)
    Qualitative Randomness: The third study provides a qualitative comparison. We analyzed the spectrograms of audio generated using both simple and random selection modes to visually illustrate the structural differences in the resulting perturbation patterns.
6. Results
This section presents the comprehensive evaluation of our proposed Spectral Masking & Interpolation Attack (SMIA). We analyze its effectiveness against state-of-the-art anti-spoofing countermeasures (CMs) and voice authentication systems (VAS) across the various scenarios outlined in our methodology.
6.1. End-to-End System Evaluation on ASVspoof 2019
The results of our evaluation on the ASVspoof 2019 dataset, a standardized benchmark for voice security research, are presented in Table 5. Crucially, our findings demonstrate that the SMIA attack is highly effective even against these layered defenses, achieving a formidable ASR of up to 99.3% against combined VAS and CM security pipelines.
A direct numerical comparison with prior works, such as Kassis et al. (Kassis and Hengartner, 2023), was not conducted, as they often report different metrics that are not directly comparable to the ASR used in our analysis. Similarly, a comparison with Hai et al. (Hai et al., 2023) is not included, as their analysis did not extend to the combined VAS and CM systems that are the focus of our evaluation. Therefore, the purpose of this experiment is not to draw a direct comparison, but to establish a robust and reproducible performance baseline for the SMIA attack. These results provide a clear benchmark for any future research conducted on this standard dataset.
Table 5. Attack Success Rate (ASR) (%) of SMIA against end-to-end systems (VAS and CMs) on the ASVspoof 2019 dataset.
VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
X-Vectors 99.3 80 99.3
DeepSpeaker 67.9 47.3 69.6
6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (Panayotov et al., 2015), Table 6 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations and never falling below a formidable 82.7% success rate. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses and underscoring the urgent need for more resilient countermeasures.
Table 6. ASR (%) of SMIA against end-to-end systems on the LibriSpeech dataset. The systems are composed of VAS and CMs.
VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
X-Vectors 100 100 100
DeepSpeaker 100 93.3 82.7
6.3. Targeted System Evaluations
To gain deeper insights into the vulnerabilities of individual system components, we analyzed the results of our targeted attacks against CMs and VAS separately.
6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (et al., 2019). The results, presented in Table 7, are unequivocal. SMIA substantially outperforms both state-of-the-art baselines across all tested countermeasures, achieving a near-perfect 99.3% ASR and demonstrating a significant advancement in attack efficacy.
Table 7. Performance comparison (ASR) of SMIA against state-of-the-art attacks on the ASVspoof 2019 dataset.
Attack Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
SMIA 99.3 81.1 99.3
SiFDetectCracker (Hai et al., 2023) 80.4 75.8 84.1
Kassis et al. (Kassis and Hengartner, 2023) 12.03 15.82
6.3.2. Attacks on Anti-Spoofing Systems using LibriSpeech
A subsequent experiment using adversarial audio from the LibriSpeech pipeline revealed catastrophic vulnerabilities across all tested CMs. As detailed in Table 8, the results were definitive: our SMIA attack achieved a perfect 100% ASR against every state-of-the-art model evaluated.
Table 8. ASR (%) of SMIA against individual anti-spoofing CMs using adversarial audio generated from the LibriSpeech pipeline.
Model ASR (%)
RawNet2 100
RawGAT-ST 100
RawPC-Darts 100
6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 9, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (Li et al., 2017), X-Vectors (Snyder et al., 2018), and the commercial Microsoft Azure SV (Microsoft, [n. d.]), peaking at a perfect 100%.
Table 9. Performance of the SMIA attack against Voice Authentication Systems (DeepSpeaker, X-Vectors, and Microsoft Azure) across SV, CSI, and OSI tasks.
Attack Task Model Recall ASR (%) F1-score EER
SMIA SV Deep Speaker 0.99 97.5 0.67 0.49
SV X-Vectors 1 100 0.67 0.5
SV Microsoft Azure 0.76 98 0.56 0.6
CSI Deep Speaker 0.96 87 0.67 0.45
CSI X-Vectors 0.98 97 0.66 0.49
OSI Deep Speaker 0.95 88 0.68 0.46
OSI X-Vectors 0.98 97 0.66 0.49
Kassis et al. (Kassis and Hengartner, 2023) SV X-Vectors - 100 - -
6.3.4. Computational Efficiency
For a holistic evaluation, we further compared the time cost of SMIA against the baselines. As summarized in Table 10, our framework consistently required fewer iterations and lower computational time to achieve successful attacks. Specifically, SMIA achieved near-perfect attack success within minutes, while Hai et al. (Hai et al., 2023) required significantly longer runtimes under comparable settings. This reduction in time cost emphasizes the practical efficiency of SMIA, making it not only more effective but also more feasible for real-world adversarial scenarios.
Table 10. Average runtime comparison of SMIA and Hai et al. (Hai et al., 2023).
Attack Avg Time (sec)
SMIA 28
Hai et al. (Hai et al., 2023) 185
6.4. Robustness in Real-World Scenarios
The results of our simulated transmission channel experiments are shown in Table 11. This table compares the attack success rate in a clean digital environment versus simulated over-the-air and over-the-line conditions. While there is a slight degradation in performance, the SMIA attack remains highly effective in both scenarios, confirming its robustness and practical applicability in real-world environments.
Table 11. Robustness of the SMIA attack against Voice Authentication Systems under simulated over-the-line and over-the-air transmission channels.
Channel Model Recall ASR(%) F1-score EER
Over-the-line Deep Speaker 0.99 99.5 0.66 0.49
Over-the-line X-Vectors 1 100 0.67 0.5
Over-the-air Deep Speaker 1 100 0.67 0.5
Over-the-air X-Vectors 1 100 0.67 0.5
6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (Hai et al., 2023), as visualized in Figure 4. The spectrograms for the Hai et al. attack (Figures 4(b) and 3(g)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
(a) Original
(b) Hai et al. (Hai et al., 2023)
(c) SMIA–Mask
(d) SMIA–Interpol.
(e) SMIA–Hybrid
(f) Original
(g) Hai et al. (Hai et al., 2023)
(h) SMIA–Mask
(i) SMIA–Interpol.
(j) SMIA–Hybrid
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (Hai et al., 2023) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (Hai et al., 2023) (b, g), which create a detectable pattern. In contrast, the perturbations from all SMIA modes (c, d, e) and (h, i, j) are more subtle and randomly distributed, making them significantly harder to detect.
7. Ablation Study
7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 12, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. The key takeaway is that the Interpolation mode proved to be the most potent overall, achieving a perfect 100% ASR in multiple configurations and indicating its superior ability to bypass certain defense combinations.
Table 12. Ablation study of SMIA attack modes. The table shows the ASR (%) for the Masking, Interpolation, and Hybrid modes against various combined CM and VAS systems.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Masking X-Vectors 78.3 68.3 93.3
DeepSpeaker 48.3 71.6 46.7
Interpolation X-Vectors 100 100 68.3
DeepSpeaker 81.3 77.8 23.3
Hybrid X-Vectors 73.3 70 71
DeepSpeaker 50 70 50
7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 13 reveal a crucial insight: while both approaches are exceptionally potent, introducing randomness provides a decisive performance advantage against certain advanced defenses. For example, against the DeepSpeaker (Li et al., 2017) and RawGAT-ST (Tak et al., 2021a) combination, the Random mode boosts the ASR from 77.8% to 93.3%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
Table 13. Ablation study on the impact of randomness. The table compares the ASR (%) of a deterministic Simple vs a stochastic (Random) perturbation strategy against various combined CM and VAS systems.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Simple X-Vectors 100 99.7 100
DeepSpeaker 96.8 77.8 82.7
Random X-Vectors 100 100 78.3
DeepSpeaker 100 93.3 73
7.3. Stealth Analysis of Simple vs. Random
The superiority of the Random mode, demonstrated by its higher ASR in Table 13, is directly linked to its enhanced stealth. Figure 4 provides a clear visual explanation for this advantage. The Simple mode attacks (top row) exhibit a fixed, deterministic perturbation pattern, creating a consistent and thus more easily detectable signature for forensic analysis. In contrast, the Random mode attacks (bottom row) show a stochastic, varied pattern that lacks predictability. The key takeaway is that this randomness makes the attack significantly harder to fingerprint, which explains its more robust performance and makes it a more effective strategy.
(a) SMIA Simple Mask
(b) SMIA Simple Interpolation
(c) SMIA Simple Hybrid
(d) SMIA Random Mask
(e) SMIA Random Interpolation
(f) SMIA Random Hybrid
Figure 4. Visual analysis of Simple vs. Random perturbation. The top row displays the fixed patterns of the Simple mode. The bottom row showcases the varied and unpredictable patterns of the Random mode, which are inherently stealthier. This visual difference in predictability helps explain the superior ASR of the Random mode shown in Table 13. The white boxes highlight the areas that exhibit differences between the modes’ spectrograms.
8. Discussion
Spectral Masking and Interpolation Attack (SMIA) represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors and commercial systems such as Microsoft Azure SV, demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat. Furthermore, the attack’s robustness in simulated over-the-air and over-the-line conditions confirms its practicality beyond purely digital environments.
Another crucial aspect of the attack is its stealth. The visual analysis of spectrograms and the ablation study on randomness reveal that the stochastic nature of the Random mode is not only more effective but also significantly harder to detect. The subtle, non-uniform perturbations are inherently more resilient to simple forensic countermeasures. The core mechanism of SMIA, which targets inaudible, low-energy frequency regions, allows it to deceive countermeasures without substantially altering the speaker’s core vocal characteristics needed for verification. This surgical approach effectively resolves the common challenge in adversarial attacks: bypassing the anti-spoofing system while preserving the biometric identity required to fool the voice authentication system.
Ultimately, this research presents a compelling argument for a fundamental rethinking of voice security. The high success rates of SMIA against state-of-the-art, layered defenses indicate that static, pattern-based detection is insufficient. As stated in the conclusion, these findings serve as an urgent call for a paradigm shift toward dynamic, context-aware security frameworks capable of co-evolving with the threat landscape.
9. Future Work
Building on the foundation of this research, several promising avenues for future work can be explored to enhance attack efficiency and, in turn, develop more robust defenses.
First, to improve the attack’s computational efficiency, the current iterative feedback loop could be replaced with a more sophisticated optimization algorithm. Techniques such as Bayesian optimization or other gradient-free methods could search the parameter space for successful perturbations, potentially reducing the number of queries and overall time required to craft an effective adversarial sample. An even more advanced approach involves training a Deep Neural Network (DNN) to act as a perturbation generator. Such a network could take the initial spoofed audio as input and directly output a perturbed sample or the parameters to generate one, replacing the iterative search with a single, rapid forward pass.
Second, to enhance the precision and stealth of the attack, a supplementary network could be developed to identify and localize the specific time-frequency bins in a synthetic audio sample that are most likely to be flagged as artificial. The SMIA framework could then focus its perturbations on these specific regions, leading to a more targeted and potentially even stealthier attack.
Finally, the insights from SMIA should be leveraged to build proactive defenses. A critical next step is to use the adversarial examples generated by this attack to adversarially train the next generation of voice authentication and anti-spoofing models. By exposing models to these advanced attacks during training, they can learn to recognize and reject such sophisticated manipulations, forming the basis of the dynamic defense frameworks this paper calls for. Further research could also explore perturbations in alternative acoustic domains, such as the cepstral domain or the raw waveform, to uncover new vulnerabilities and inform the design of more comprehensive, multi-domain defense strategies.
10. Conclusion
This study introduces the Spectral Masking and Interpolation Attack (SMIA), a novel black-box method that reveals critical vulnerabilities in modern voice security infrastructure. By strategically manipulating imperceptible frequency regions to naturalize synthetic audio, SMIA crafts adversarial samples that sound authentic yet consistently bypass defenses, achieving a 100% ASR against standalone countermeasures, 97.5% against voice authentication systems, and a strong 82% against combined VAS/CM pipelines. These findings conclusively demonstrate that current security postures, including layered defenses, are insufficient against adaptive threats. This research serves as an urgent call for a paradigm shift away from static models and toward dynamic, context-aware security frameworks capable of evolving with the threat landscape to protect the future of voice biometrics.
Acknowledgement
This research is supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4003.