### v4

# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

Abstract.
Voice Authentication Systems (VAS) are increasingly integrated into high-security sectors such as banking and healthcare. Despite their improvements in the previous years, they still face severe vulnerabilities from sophisticated threats, such as deepfakes and adversarial attacks. The emergence of realistic voice cloning complicates detection, as systems struggle to distinguish authentic from synthetic audio. While anti-spoofing countermeasures (CMs) exist to mitigate these risks, many rely on static detection models that can be bypassed by novel adversarial methods, leaving a critical security gap. To demonstrate this vulnerability, we propose the Spectral Masking and Interpolation Attack (SMIA), a novel method that strategically manipulates inaudible frequency regions of AI-generated audio. By altering the voice in imperceptible zones to the human ear, SMIA creates adversarial samples that sound authentic while deceiving CMs. We conducted a comprehensive evaluation of our attack against state-of-the-art (SOTA) models across multiple tasks, under simulated real-world conditions. SMIA achieved SOTA attack results, demonstrating a high success rate in bypassing combined VAS/CM systems, standalone speaker verification systems, and dedicated countermeasures. These findings conclusively show that current security postures are insufficient against adaptive adversarial attacks, highlighting the urgent need for next-generation, dynamic defenses.
Machine Learning Security, Voice Authentication, Anti-Spoofing, Voice Cloning, Adversarial Spoofing Attacks
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM Asia Conference on Computer and Communications Security; June 01–05, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/2026/06ccs: Security and privacy Biometricsccs: Computing methodologies Neural networks
1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (Singh et al., 2012; Algabri et al., 2017; Rose, 2006). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (Renz et al., 2023; Springer, 2023; Techlocity, 2021; Khan and Zahid, 2010).
Recent advancements in deep learning have revolutionized the Voice Authentication System (VAS), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (Lan et al., 2022).
In parallel, deep learning has transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (Liao et al., 2024; Qin et al., 2023; Amazon Web Services, 2025) and voice conversion (VC) (Li et al., 2023; Huang et al., 2024) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass VAS (Shonesy, 2015; Alali and Theodorakopoulos, 2025).
To combat these deepfake attacks, various countermeasures (CMs), like RawNet2 (Tak et al., 2021b), RawGAT-ST (Tak et al., 2021a), and RawPC-Darts (Ge et al., 2021), have been developed to distinguish between genuine and synthetic speech. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: bypassing the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (Hai et al., 2023; Kassis and Hengartner, 2023).
A compelling example is SiFDetectCracker (Hai et al., 2023), an adversarial attack that can deceive SOTA CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are effective in real-world scenarios.
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. Kassis et al. (Kassis and Hengartner, 2023) introduced a model-agnostic attack that bypasses these CMs without relying on query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (Amazon, 2025) and RawPC-Darts (Ge et al., 2021). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of SOTA, hardened VA systems.
Our main contributions are summarized as follows:
  • We introduce the Spectral Masking and Interpolation Attack (SMIA). This novel black-box adversarial spoofing attack exploits inaudible regions of AI-generated audio to bypass both VAS and CMs.
  • We conduct extensive evaluations of SMIA against widely adopted SOTA open-source and commercial VAS, and SOTA CMs.
  • Our results demonstrate that SMIA significantly outperforms current state-of-the-art attacks by addressing critical methodological gaps, thereby creating a more robust and effective threat.
Unlike prior work, which explored TTS-based attacks targeted at VAS and CMs in isolation or relied on transferability-based attacks with limited effectiveness, we present a black-box attack that simultaneously compromises both. Our evaluations span a wide range of models, tasks, and real-world scenarios, demonstrating that VAS are significantly more fragile than previously assumed, underscoring the urgent need for stronger defenses and providing a solid foundation for future research in biometric security.
This paper is organized as follows. We first review the background and related work (Section 2). We then detail our SMIA methodology (Section 3) and the experimental setup (Section 4). Section 5 presents the evaluation of our attack. Section 6 reports the results of our experiments. Section 7 provides an ablation study, analyzing the impact of different components and design choices. Finally, we discuss our findings (Section 8), outline future work (Section 9), and conclude (Section 10).
2. Background and Related Works
2.1. Speaker Recognition Systems (SRS)
SRS identifies individuals through their vocal features. The authentication process typically unfolds in three key phases, as shown in Figure 1:
  1. (1)
    First, during the training phase, the system learns voice characteristics from a large corpus to develop robust voice authentication models.
  2. (2)
    Next, in the enrollment phase, a user’s voice is recorded as they speak specific phrases, and the extracted embeddings are stored.
  3. (3)
    Finally, during the recognition phase, the system compares embeddings from the newly captured voice sample to the enrolled embeddings to confirm identity.
Figure 1. Speaker Recognition System Overview. 1) Training phase. 2) Enrollment phase. 3) Verification phase.
2.2. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (Patel and Patil, 2015; De Leon et al., 2017). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has shifted towards deep learning methods (Lavrentyeva et al., 2019), with SOTA now driven by Self-Supervised Learning (SSL) models like Wav2Vec2.0 (Baevski et al., 2020) and WavLM (Chen et al., 2022). WavLM has shown superior performance by capturing fine-grained artifacts commonly found in deepfakes. However, pre-training these large-scale SSL models takes a lot of time and computational resources. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (Lavrentyeva et al., 2017) or the AASIST model (Jung et al., 2022), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (Tak et al., 2021b) demonstrated strong performance in the ASVspoof 2019 challenge (et al., 2019), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as SOTA alternatives. RawPC-Darts (Ge et al., 2021) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, RawGAT-ST (Tak et al., 2021a) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
2.3. Attacks Taxonomies
Attacks are categorized based on the attacker’s goal to:
  • Targeted Attacks: Forcing the system to recognize the input as a specific enrolled speaker.
  • Untargeted Attacks: Causing misclassification as any incorrect speaker without control over the output identity.
These attacks can also be categorized based on the attacker’s knowledge:
  • White-box Attack: The attacker fully knows the system, including its architecture, parameters, training data, and feature extraction methods. This enables precise gradient-based crafting of adversarial examples.
  • Grey-box Attack: The attacker possesses partial knowledge, such as the feature representation or model architecture, but lacks access to full internals like trained weights. Attacks typically exploit transferability or inferred system behavior.
  • Black-box Attack: The attacker lacks any internal knowledge and treats the system like an oracle. They rely on input-output queries. Attack methods usually involve optimizing queries or using transfer attacks with surrogate models.
Table 1. Summary of Key Deepfake and Adversarial Spoofing Attacks on Speaker Recognition Systems
Paper Year Tasks Attack Type System Target
Deepfake Voice Attacks
Saxena et al. (Shonesy, 2015) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (Alali and Theodorakopoulos, 2025) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
Adversarial Spoofing Attacks
Liu et al. (Liu et al., 2019) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (Gomez-Alanis et al., 2022b) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (Hai et al., 2023) 2023 CM Black-box CM
Kassis et al. (Kassis and Hengartner, 2023) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
2.4. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (Ren et al., 2019, 2020; Kim et al., 2020; Kaneko and Kameoka, 2018; Kameoka et al., 2018; Qian et al., 2019; Li et al., 2023; Qin et al., 2023; Wang et al., 2024). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (Almutairi and Elgibreen, 2022).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (Shonesy, 2015). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: SOTA SRS failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (Alali and Theodorakopoulos, 2025) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and SRS remain vulnerable when exposed to partial speech generated through zero-shot TTS and VC models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
2.5. Adversarial Spoofing Attacks
These attacks typically involve adding subtle perturbations to AI-generated audio, crafted by speech synthesis models, to fool CMs or both the VAS and CMs.
2.5.1. White-box Attacks
Early work by Liu et al. (Liu et al., 2019) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to CM systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51% (absolute) across ASVspoof 2019 logical and physical access protocols (et al., 2019).
In a related study, Alanis et al. (Gomez-Alanis et al., 2022b) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.
2.5.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. (Hai et al., 2023) proposed SiFDetectCracker, a framework targeting CMs. Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), RawPC-Darts (Ge et al., 2021), and Deep4SNet (Ballesteros et al., 2021), using the ASVspoof 2019 dataset (et al., 2019). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (Kassis and Hengartner, 2023) presents a novel methodology for bypassing VAS, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (Ge et al., 2021) and RawGAT-ST (Tak et al., 2021a), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
3. Methodology
Figure 2. Our Spectral Masking & Interpolation Attack (SMIA) that targets both Voice Authentication and Anti-Spoofing Systems.
A combined voice authentication and anti-spoofing system processes a given voice sample $x\in\mathbb{R}^{m}$ to perform two key tasks: SV and CM. The system grants access only when the sample is verified as belonging to the claimed identity and determined to be a genuine, live utterance.
This integrated system can be formally defined with a speaker verification model $g(x,e)$ and an anti-spoofing model $f(x)$. For a given input $x$ and an enrolled embeddings $e$, the system’s decision $S(x,e)$ is given by:
$S(x,e)=\begin{cases}\text{accept}&\text{if }g(x,e)=\text{accept}\land f(x)=\text{real}\\ \text{reject}&\text{otherwise}\end{cases}$
The objective of our attack is to generate an adversarial sample $x^{\prime}$ from an initial spoofed utterance $x$ that bypasses this combined system. This requires crafting an $x^{\prime}$ that is both accepted by the speaker verification model and misclassified as real by the anti-spoofing model. In short, our goal is to ensure $S(x^{\prime},e)=\text{accept}$.
In this section, we first present Spectral Masking and Interpolation Attack (SMIA), which manipulates these features to bypass the countermeasure while preserving the speaker characteristics needed to fool the verification model, before delving into its details.
3.1. Threat Model
Our threat model considers an attacker whose primary goal is to gain unauthorized access to a victim’s account on an SV system by mimicking their voice. The success of the attack is defined by a single successful authentication event.
The attacker operates under a black-box assumption, meaning they have no internal knowledge of the system’s architecture, including the specific VAS and CMs employed. The attacker’s capabilities are as follows:
  • Data Collection: They can collect voice samples of the victim from publicly available sources (e.g., social media), requiring a minimum of 10 seconds of audio to initiate the attack.
  • Sample Generation: They can generate a spoofed audio sample using any voice-cloning tool (e.g., Fish Speech APIs (Liao et al., 2024) or methods used in ASVspoof 2019 (et al., 2019)).
  • Iterative Optimization: The attacker can repeatedly submit generated audio to the authentication endpoint to observe feedback. They use this feedback to iteratively optimize the sample to bypass defenses.
Regarding the feedback loop, we consider two specific system response scenarios:
  • Score and Label: The system returns both a continuous confidence score and the final decision (like Azure SV (Microsoft, )).
  • Label Only: The system returns only a binary pass/fail decision.
Our method is effective in both scenarios. The attacker’s actions are considered stealthy, as a successful login does not interfere with or alert the legitimate user. Also, the attack is applicable to both text-dependent and text-independent SV systems.
3.2. Overview
We propose a novel black-box adversarial attack framework designed to circumvent both VAS and their corresponding CMs. The attack operates by combining high-fidelity voice synthesis with an adaptive, iterative perturbation process.
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—the attacker can employ Fish Speech APIs (Liao et al., 2024), ASVSpoof 2019 challenge tools (et al., 2019), or any modern voice cloning system.
Second, this synthesized audio serves as the input to our core attack mechanism—the Spectral Masking and Interpolation Attack (SMIA) system. SMIA uses a black-box optimization algorithm to modify the audio using our proposed Spectral Masking and Interpolation technique. The resulting adversarial audio is submitted to the target system (VAS and CM), and the prediction outcome is used as feedback to update the SMIA parameters. This process is repeated, progressively refining the audio until it successfully deceives the target model. Our attack framework is thus composed of two integral components:
  • Black-Box Optimization Framework: A feedback-driven loop that queries the target model and uses its decision outcome to adaptively update the params to guide the generation of the adversarial audio.
  • Spectral Masking and Interpolation: A stealthy perturbation method that introduces distortions by targeting perceptually insignificant spectral regions, thereby preserving the audio’s naturalness.
Together, these components create a robust and adaptive feedback-driven process that dynamically refines its adversarial strategy, making the attack highly effective and difficult to defend against.
3.3. Black-Box Optimization (BBO) Framework
This component serves as the “engine” of the attack, systematically searching for the optimal parameters for the spectral perturbation module (detailed in Section 3.4). As defined in our black-box threat model (Section 3.1), we assume no gradient information or internal model access is available.
To efficiently navigate the search space, we employ a Tree-structured Parzen Estimator (TPE). TPE is a Bayesian optimization technique ideally suited for this task, as it builds probabilistic models of “good” ($l(\theta)$) and “bad” ($g(\theta)$) parameters to suggest candidates that maximize the likelihood of success. Our framework is designed to be flexible, adapting its objective function based on the feedback tuple $(s,l)$ returned directly by the target model, where $s$ is the confidence score and $l$ is the decision label.
Unlike varying threshold-based approaches, our optimization utilizes the target’s output directly. The optimization objective $J(\theta,m)$ is defined adaptively:
[…]
In the primary scenario where the target returns a continuous score $s$, the TPE optimizer uses this value directly as the scoring function to guide the search toward higher confidence regions. In strict decision-based scenarios, the score collapses to a binary value. Crucially, the returned label $l$ serves as the termination criterion for the attack loop; the optimization halts immediately once a successful label is received, regardless of the score magnitude.
The TPE optimizer models the parameter distributions as shown in Equations 3 and 4, where $J^{*}$ is a quantile of observed objective scores. It suggests new candidates that maximize the ratio $l(\theta)/g(\theta)$, efficiently focusing the search on promising regions.
(3) $l(\theta)$ $=p(\theta|J(\theta,m)>J^{*})$
(4) $g(\theta)$ $=p(\theta|J(\theta,m)\leq J^{*})$
To define the optimizer’s constraints, we bound its search space and budget. The hyperparameter ranges are: the threshold $t_{db}$ is searched between -60.0 dB and -10.0 dB; the probability mean $\mu$ is searched between 0.0 and 1.0; and the probability standard deviation $\sigma_{p}$ is searched between 0.0 and 0.75. The framework performs this for a maximum of $N_{iter}=100$ trials, as detailed in Algorithm 1.
Algorithm 1 Black-Box Optimization Attack
1:Input: Enrolled user ID $U_{id}$, Audio Verification Sample $A_{V}$, Max iterations $N_{iter}$
2:Output: Attack success status and best result $R_{best}$
3:procedure Attack($U_{id}$, $A_{V},N_{iter}$)
4:  $M\leftarrow\{\text{`interpolate', `mask', `hybrid'}\}$ $\triangleright$ Attack modes
5:  $S_{best}\leftarrow$-$\infty$
6:  $R_{best}\leftarrow\text{None}$
7:  $\text{success}\leftarrow\text{False}$
8:  for $m\in M$ do
9:   Initialize BBO sampler $\mathcal{S}_{m}$ $\triangleright$ $\theta=\{t_{db},\mu,\sigma_{p}\}$
10:   for $i\leftarrow 1\text{ to }N_{iter}$ do
11:     $\theta\leftarrow\mathcal{S}_{m}.\text{suggest\_parameters}()$
12:     $A_{C}\leftarrow\textbf{ApplyPerturbations}(A_{V},\theta,m)$
13:     $(s,l)\leftarrow\text{Authenticate}(\textbf{$U_{id}$},A_{C})$
14:     $\mathcal{S}_{m}.\text{report\_result}(s,\theta)$
15:     if $s>S_{best}$ then
16:      $S_{best}\leftarrow s$
17:      $R_{best}\leftarrow(s,l,\theta,m)$
18:     end if
19:     if $l=\text{`success'}$ then
20:      $\text{success}\leftarrow\text{True}$
21:      break
22:     end if
23:   end for
24:   if $\text{success}=\text{True}$ then
25:     break
26:   end if
27:  end for
28:  return $R_{best}$
29:end procedure
3.4. Spectral Masking and Interpolation
At the core of our attack is a perturbation module, ApplyPerturbation. The function of this module is to introduce subtle, controlled distortions into an audio signal’s spectrogram. By design, it operates exclusively on low-energy (quiet) time-frequency bins, which are perceptually less significant to the ear.
This strategic targeting is controlled by the three key parameters $\theta=(t_{db},\mu,\sigma_{p})$ from our optimizer (as shown in Algorithm LABEL:alg:spectral_perturbation):
  • $t_{\text{db}}$: A decibel threshold used to identify the quiet parts of the spectrogram.
  • $\mu$: The mean probability for selecting a quiet bin.
  • $\sigma_{p}$: The standard deviation of the probability, introducing randomness to the selection.
Our algorithm uses $\mu$ and $\sigma_{p}$ to define the probability $p$ of a quiet bin being perturbed. This value $p$ is sampled from a Normal distribution $N(\mu,\sigma_{p}^{2})$ and then clipped to the valid range $[0,1]$. This stochastic approach (sampling $p$ for each attack) prevents a fixed, detectable perturbation pattern.
This overall method ensures that the resulting adversarial audio remains natural-sounding and free of obvious artifacts. The module employs three primary techniques to manipulate the audio’s underlying acoustic features: masking, interpolation, and a hybrid approach.
Algorithm 2 ApplyPerturbations: Spectral Perturbation Function
1:Audio signal $x$, parameter vector $\theta$, attack mode modemode
2:Adversarial audio $x_{adv}$
[…]
4:procedure SampleProbability($\mu,\sigma_{p}$)
5:  $p_{raw}\leftarrow\text{SampleFromNormal}(\mu,\sigma_{p})$
6:  return $\text{Clip}(p_{raw},0.0,1.0)$
7:end procedure
8:procedure ApplyPerturbations($x,\theta,mode$)
9:  $\text{mag},\text{phase}\leftarrow\text{STFT}(x)$ $\triangleright$ STFT: Short-Time Fourier Transform that converts $x$ into time–frequency magnitude and phase
10:  $\text{quiet\_mask}\leftarrow\text{mag}<t_{db}$
11:  $p\leftarrow\text{SampleProbability}(\mu,\sigma_{p})$
12:  $\text{random\_mask}\leftarrow\text{UniformRandom}(\text{shape}(\text{mag}))<p$
13:  $\text{final\_mask}\leftarrow\text{quiet\_mask}\land\text{random\_mask}$
14:  $\text{mag\_out}\leftarrow\text{mag}$
15:  if $mode=\text{`interpolate'}$ or $mode=\text{`hybrid'}$ then
16:   $\text{mag\_out}\leftarrow\text{Interpolate}(\text{mag\_out},\text{final\_mask})$
17:  end if
18:  if $mode=\text{`mask'}$ or $mode=\text{`hybrid'}$ then
19:   $\text{mag\_out}[\text{final\_mask}]\leftarrow 0$
20:  end if
21:  $x_{adv}\leftarrow\text{ISTFT}(\text{mag\_out},\text{phase})$ $\triangleright$ ISTFT reconstructs the time-domain signal from modified spectral components
22:  return $x_{adv}$
23:end procedure
3.4.1. Perturbation Modes
Masking
In this mode, the module takes a straightforward approach: it resets the magnitude of the selected quiet and randomized time-frequency bins to zero. This effectively silences these specific parts of the signal, creating a subtle distortion by removing minimal acoustic information.
Interpolation
This mode aims to replace the targeted quiet bins with new values that are consistent with the surrounding stable, non-quiet parts of the signal. For each frequency channel in the spectrogram, the module performs the following steps:
  • It identifies all reliable, high-energy (non-quiet) points over time, using them as anchors.
  • It constructs a one-dimensional linear interpolation function based on the time indices and magnitude values of these anchor points.
  • It uses this function to generate a new temporal contour for the entire frequency bin, and then selectively replaces the original magnitudes only at the targeted quiet locations with the corresponding values from this new contour.
Essentially, it reconstructs the quiet segments based on the audio’s natural temporal contour within that specific frequency. The magnitude $y$ at a time $x$ falling between two known anchor points $(x_{0},y_{0})$ and $(x_{1},y_{1})$—where $x$ represents time and $y$ represents magnitude—is calculated using the standard linear interpolation formula:
(5) $y(x)=y_{0}+(x-x_{0})\frac{y_{1}-y_{0}}{x_{1}-x_{0}}$
This process ensures that the filled-in values are temporally smooth and plausible, making the alteration difficult to perceive.
Hybrid
The hybrid mode offers the most flexibility by combining both masking and interpolation. It applies a mix of both techniques to the targeted bins, creating a more complex and varied spectral perturbation. This allows the attack to leverage the distinct advantages of each method simultaneously.
4. Experimental Setup
Our experimental setup leverages the LibriSpeech (Panayotov et al., 2015) and ASVSpoof2019 (et al., 2019) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), and RawPC-Darts(Ge et al., 2021). For the voice authentication stage, we employ three distinct systems: X-Vectors (Snyder et al., 2018), DeepSpeaker (Li et al., 2017), and the commercial Azure Speaker Verification API (Microsoft, ). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations:
  • Combined systems integrating both anti-spoofing and voice authentication.
  • Anti-spoofing systems in isolation.
  • Voice Authentication Systems in isolation.
4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (et al., 2019) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (Panayotov et al., 2015) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
4.1.1. ASVSpoof 2019
The ASVSpoof 2019 dataset is the primary benchmark for evaluating spoofing countermeasures. We use the Logical Access (LA) Evaluation partition, whose partitions are statistically summarized in Table 2.
Table 2. Statistics of the ASVSpoof 2019 LA Dataset and Experimental Subset.
Partition Speakers (M/F) Total Speakers Total Utt.
Training 8 / 12 20 25,380
Development 8 / 12 20 24,844
Evaluation 29 / 39 68 71,237
4.1.2. LibriSpeech
We used the test-clean subset of the LibriSpeech ASR corpus. This subset provides high-quality audio from 40 unique speakers (20 male and 20 female). To simulate a realistic user verification scenario, we selected a specific number of samples from this subset. From each of the 40 speakers, we designated 5 utterances for enrollment to create their voiceprint and a separate set of 5 utterances for verification trials. This resulted in a total of 400 selected audio files (10 per speaker) for our evaluation.
4.2. Authentication Models
To ensure a comprehensive evaluation, we test our attack against a diverse set of three SOTA SV systems. This set includes two well-known open-source models and one commercial black-box API, allowing us to assess performance on both accessible academic benchmarks and proprietary, real-world systems.
  • X-Vectors (Snyder et al., 2018): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (Ravanelli et al., 2021) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (Li et al., 2017) [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (Microsoft, ) [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (et al., 2019). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (Tak et al., 2021b) [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (et al., 2019). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (Tak et al., 2021a) [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (et al., 2019).
  • RawPC-DARTS (Ge et al., 2021) [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (et al., 2019), making it a critical system to evaluate against.
5. Evaluation Methodology
To comprehensively assess the effectiveness of our proposed attack, we designed a series of rigorous experiments targeting SOTA VAS and CMs. Our methodology is structured to benchmark performance against prior work, evaluate end-to-end systems, and test robustness in realistic scenarios.
  1. (1)
    End-to-End System Evaluation on ASVspoof: To create a robust evaluation subset, we randomly selected 30 speakers from the ASVspoof 2019 LA Evaluation partition (et al., 2019). For each selected speaker, we used 5 of their genuine bona fide audio samples for enrollment in the VAS. We then randomly selected 5 of their corresponding pre-existing spoofed audio files to serve as the base samples for our SMIA attack, resulting in 150 attack trials (30 speakers $\times$ 5 samples).
  2. (2)
    End-to-End System Evaluation on LibriSpeech: Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (Panayotov et al., 2015). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (Liao et al., 2024). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
  3. (3)
    Targeted System Evaluations: To gain deeper insights into the vulnerabilities of individual system components, we conduct targeted experiments against CMs and VAS separately.
    1. (a)
      Benchmark Attacks on Anti-Spoofing Systems (ASVspoof): To ensure a fair comparison with prior work (Kassis and Hengartner, 2023; Hai et al., 2023), we applied our SMIA framework to spoofed audios from the ASVspoof 2019 dataset (et al., 2019). This methodology allows for a direct performance benchmark against existing methods. Specifically, we targeted all 13 logical access spoofing attack types (A07-A19) present in the evaluation partition. For each attack type, we randomly selected 50 audio samples, creating a comprehensive test set of 650 spoofed utterances (13 types $\times$ 50 samples) to ensure our attack’s generalizability.
    2. (b)
      Attacks on Anti-Spoofing Systems (LibriSpeech): We conducted a dedicated experiment targeting CMs. Test audios were generated using LibriSpeech (Panayotov et al., 2015) and Fish Speech APIs (Liao et al., 2024), then processed by our SMIA attack to test the vulnerability of various CMs.
    3. (c)
      Attacks on Voice Authentication Systems (VAS): We extensively evaluated VAS vulnerabilities using audio generated from LibriSpeech (Panayotov et al., 2015) and Fish Speech (Liao et al., 2024). The attacks were tested against VAS models across three biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
    4. (d)
      Computational Efficiency Evaluation: We evaluated SMIA’s efficiency against SiFDetectCracker (Hai et al., 2023), measuring the iterations and time required for a successful attack.
    5. (e)
      Robustness in Real-World Scenarios: To assess practical applicability, we tested the generated audio against VAS systems in two simulated real-world environments: over-the-air (speaker-to-microphone) and over-the-line (VoIP codec simulation).
  4. (4)
    Stealth and Detectability Analysis: To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative visual analysis of the audio spectrograms. We compare the adversarial samples generated by SMIA against those from SiFDetectCracker (Hai et al., 2023) to show that our method avoids the structured, conspicuous artifacts that create a detectable pattern. This highlights our method’s superior stealth and resilience against basic forensic analysis.
  5. (5)
    Ablation Study: To understand which parts of our attack contribute most to its success, we conducted these ablation studies:
    1. (a)
      Attack Modes: The first study focuses on our perturbation strategies. We systematically compared the performance of the three distinct modes—Masking, Interpolation, and Hybrid—to determine which method is most effective at bypassing the target systems.
    2. (b)
      Randomness: The second study investigates the importance of our stochastic approach. We evaluated the attack’s performance both with and without the random selection of quiet bins. This analysis clarifies how randomness contributes to the attack’s flexibility and ability to evade detection.
    3. (c)
      Qualitative Randomness: The third study provides a qualitative comparison. We analyzed the spectrograms of audio generated using both simple and random selection modes to visually illustrate the structural differences in the resulting perturbation patterns.
6. Results
This section presents the comprehensive evaluation of our proposed SMIA. We analyze its effectiveness against SOTA CMs and VAS across the various scenarios outlined in our methodology.
6.1. End-to-End System Evaluation on ASVspoof
The results of our evaluation on the ASVspoof 2019 dataset, a standardized benchmark for voice security research, are presented in Table 3. Crucially, our findings demonstrate that the SMIA attack is highly effective even against these layered defenses, achieving a formidable ASR of up to 97% against combined VAS and CM security pipelines.
Our evaluation shows that the proposed method significantly outperforms Kassis et al. (Kassis and Hengartner, 2023). In contrast, a direct comparison with Hai et al. (Hai et al., 2023) remains unfeasible, as their work does not address the combined VAS and CM systems utilized in our analysis. Consequently, this experiment provides evidence of superior performance where comparison is possible, while also providing a clear benchmark for future research on this specific dataset configuration.
Table 3. ASR (%) of SMIA against end-to-end systems (VAS and CMs) on the ASVspoof 2019 dataset.
Group VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-Darts
Ours DeepSpeaker 59 71 30
X-Vectors 97 93 82
Kassis et al. X-Vectors - 14.25 9.67
6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (Panayotov et al., 2015), Table 4 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses.
However, a key finding in Table 4 is the notable ASR drop to 66.5% when targeting the DeepSpeaker + RawPC-Darts combination. This is not a failure of the attack but rather a critical insight into the core tradeoff of our problem. It highlights a fundamental conflict: the perturbations required to fool the robust RawPC-Darts (CM) are strong enough to degrade the speaker’s biometric identity, causing DeepSpeaker (VAS) to fail the verification.
Table 4. ASR (%) of SMIA against end-to-end systems on the LibriSpeech dataset. The systems are composed of VAS and CMs.
VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-Darts
X-Vectors 100 100 92.5
DeepSpeaker 93.5 93.5 66.5
6.3. Targeted System Evaluations
To gain deeper insights into the vulnerabilities of individual system components, we analyzed the results of our targeted attacks against CMs and VAS separately.
6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (et al., 2019). The results, presented in Table 5, are unequivocal. SMIA substantially outperforms both SOTA baselines across all tested countermeasures, demonstrating a significant advancement in attack efficacy.
Table 5. Performance comparison (ASR) of SMIA against SOTA attacks on the ASVspoof 2019 dataset.
Attack Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-Darts
SMIA 96.3 95 87
SiFDetectCracker (Hai et al., 2023) 80.4 75.8 84.1
Kassis et al. (Kassis and Hengartner, 2023) 12.03 15.82
6.3.2. Attacks on Anti-Spoofing Systems using LibriSpeech
A subsequent experiment using adversarial audio from the LibriSpeech pipeline demonstrated the effectiveness of the attack across different architectures. As detailed in Table 6, we observed a perfect 100% ASR against two of the SOTA models. For the RawDarts-PC, while not total, the attack successfully bypassed the system with an 87.9% ASR, confirming robust performance across the evaluated set.
Table 6. ASR (%) of SMIA against individual anti-spoofing CMs using adversarial audio generated from the LibriSpeech pipeline.
Model ASR (%)
RawNet2 100
RawGAT-ST 100
RawPC-Darts 87.9
6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 7, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (Li et al., 2017), X-Vectors (Snyder et al., 2018), and the commercial Microsoft Azure SV (Microsoft, ), peaking at a perfect 100%.
Table 7. Performance of the SMIA attack against Voice Authentication Systems (DeepSpeaker, X-Vectors, and Microsoft Azure) across SV, CSI, and OSI tasks.
Attack Task Model Recall ASR F1-score EER
SMIA SV Deep Speaker 0.99 97.5 0.67 0.49
SV X-Vectors 1 100 0.67 0.5
SV Microsoft Azure 0.76 98 0.56 0.6
CSI Deep Speaker 0.96 87 0.67 0.45
CSI X-Vectors 0.98 97 0.66 0.49
OSI Deep Speaker 0.95 88 0.68 0.46
OSI X-Vectors 0.98 97 0.66 0.49
Kassis et al. (Kassis and Hengartner, 2023) SV X-Vectors - 100 - -
6.3.4. Computational Efficiency
For a holistic evaluation, we compared the efficiency of SMIA against the SiFDetectCracker baseline (Hai et al., 2023) when targeting standalone CMs. We evaluated efficiency using two key metrics: (1) the average number of model queries required for a successful attack, and (2) the average computational time. The query count is a more robust, hardware-independent metric for BBO attacks, while the time cost demonstrates practical feasibility.
As summarized in Table 8, our framework demonstrates superior computational efficiency. We achieved a slight reduction in the average query count—21.6 compared to 22.3 for the baseline. However, the most significant advantage lies in the processing speed; our attack was 6.6 times faster in terms of raw time, emphasizing its practical feasibility.
Table 8. Efficiency comparison (Avg. Queries and Time) of SMIA and SiFDetectCracker (Hai et al., 2023) against standalone CMs.
Attack Avg. Queries Avg. Time (sec)
SMIA 21.6 28
SiFDetectCracker (Hai et al., 2023) 22.3 185
(a) Original
(b) Hai et al. (Hai et al., 2023)
(c) SMIA
(d) Original
(e) Hai et al. (Hai et al., 2023)
(f) SMIA
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (Hai et al., 2023) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (Hai et al., 2023) (3(b), 3(e)), which create a detectable pattern. In contrast, the perturbations from SMIA (3(c), 3(f)) are more subtle and randomly distributed, making them significantly harder to detect.
6.4. Robustness in Real-World Scenarios
The results of our simulated transmission channel experiments are shown in Table 9. This table compares the attack success rate in a clean digital environment versus simulated over-the-air and over-the-line conditions. While there is a slight degradation in performance, the SMIA attack remains highly effective in both scenarios, confirming its robustness and practical applicability in real-world environments.
Table 9. Robustness of the SMIA attack against Voice Authentication Systems under simulated over-the-line and over-the-air transmission channels.
Channel Model Recall ASR(%) F1-score EER
Over-the-line Deep Speaker 0.99 99.5 0.66 0.49
Over-the-line X-Vectors 1 100 0.67 0.5
Over-the-air Deep Speaker 1 100 0.67 0.5
Over-the-air X-Vectors 1 100 0.67 0.5
6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (Hai et al., 2023), as visualized in Figure 3. The spectrograms for the Hai et al. attack (Figures 3(b) and 3(e)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
7. Ablation Study
(a) SMIA Simple Mask
(b) SMIA Simple Interpolation
(c) SMIA Simple Hybrid
(d) SMIA Random Mask
(e) SMIA Random Interpolation
(f) SMIA Random Hybrid
Figure 4. Visual analysis of Simple vs. Random perturbation. The top row displays the fixed patterns of the Simple mode. The bottom row showcases the varied and unpredictable patterns of the Random mode, which are inherently stealthier. This visual difference in predictability helps explain the superior ASR of the Random mode shown in Table 11. The white boxes highlight the areas that exhibit differences between the modes’ spectrograms.
7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 10, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. While the Interpolation mode appears dominant against RawNet2 (Tak et al., 2021b) and RawGAT-ST(Tak et al., 2021a) (achieving 100% ASR with X-Vectors), it exhibits a critical vulnerability against RawPC-Darts (Ge et al., 2021), where its performance collapses to just 5%. Conversely, the Hybrid mode proves to be the most resilient overall. It sacrifices marginal performance on the simpler defenses to achieve a decisive 84.4% ASR against RawPC-Darts (Ge et al., 2021), demonstrating that combining strategies is essential for bypassing the most robust defense combinations.
Table 10. Ablation study of SMIA attack modes. The table shows the ASR (%) for the Masking, Interpolation, and Hybrid modes against various combined CM and VAS systems.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-Darts
Masking X-Vectors 87 75 78
DeepSpeaker 76 72 54
Interpolation X-Vectors 100 100 5
DeepSpeaker 97.7 88 2
Hybrid X-Vectors 94.4 71.1 84.4
DeepSpeaker 73.3 72.2 51.1
7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 11 reveal a crucial insight: while the deterministic approach is highly potent against most systems, introducing randomness provides a decisive performance advantage against specific advanced defenses. Notably, against the DeepSpeaker (Li et al., 2017) and RawPC-DARTS combination, the Simple mode is effectively blocked (14% ASR). However, the Random mode recovers significant capability, boosting the ASR from 14% to 66.5%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
Table 11. Ablation study on the impact of randomness. The table compares the ASR (%) of a deterministic Simple vs a stochastic (Random) perturbation strategy against various combined CM and VAS systems.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Simple X-Vectors 100 100 100
DeepSpeaker 93.6 94.4 14
Random X-Vectors 100 100 92.5
DeepSpeaker 93.5 93.5 66.5
7.3. Stealth Analysis of Simple vs. Random
The superiority of the Random mode, demonstrated by its higher ASR in Table 11, is directly linked to its enhanced stealth. Figure 4 provides a clear visual explanation for this advantage. The Simple mode attacks (top row) exhibit a fixed, deterministic perturbation pattern, creating a consistent and thus more easily detectable signature for forensic analysis. In contrast, the Random mode attacks (bottom row) show a stochastic, varied pattern that lacks predictability. The key takeaway is that this randomness makes the attack significantly harder to fingerprint, which explains its more robust performance and makes it a more effective strategy.
8. Discussion
SMIA represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors (Snyder et al., 2018) and commercial systems such as Microsoft Azure SV (Microsoft, ), demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat.
Another crucial aspect of the attack is its stealth. The visual analysis of spectrograms and the ablation study on randomness reveal that the stochastic nature of the Random mode is not only more effective but also significantly harder to detect. The subtle, non-uniform perturbations are inherently more resilient to simple forensic countermeasures. The core mechanism of SMIA, which targets inaudible, low-energy frequency regions, allows it to deceive countermeasures without substantially altering the speaker’s core vocal characteristics needed for verification. This surgical approach effectively resolves the common challenge in adversarial attacks: bypassing the anti-spoofing system while preserving the biometric identity required to fool the voice authentication system.
Ultimately, this research presents a compelling argument for a fundamental rethinking of voice security. The high success rates of SMIA against state-of-the-art, layered defenses indicate that static, pattern-based detection is insufficient. As stated in the conclusion, these findings serve as an urgent call for a paradigm shift toward dynamic, context-aware security frameworks capable of co-evolving with the threat landscape.
9. Future Work
Building on the foundation of this research, several promising avenues for future work can be explored to enhance attack efficiency and, in turn, develop more robust defenses.
First, to further enhance the attack’s computational efficiency, the optimization process itself could be replaced entirely. Our Bayesian Optimization can be replaced by a DNN to act as a perturbation generator. Such a network could take the initial spoofed audio as input and directly output a perturbed sample or the parameters to generate one, replacing the multi-query optimization search with a single, rapid forward pass.
Second, to enhance the precision and stealth of the attack, a supplementary network could be developed to identify and localize the specific time-frequency bins in a synthetic audio sample that are most likely to be flagged as artificial. The SMIA framework could then focus its perturbations on these specific regions, leading to a more targeted and potentially even stealthier attack.
Finally, the insights from SMIA should be leveraged to build proactive defenses. A critical next step is to use the adversarial examples generated by this attack to adversarially train the next generation of voice authentication and anti-spoofing models. By exposing models to these advanced attacks during training, they can learn to recognize and reject such sophisticated manipulations, forming the basis of the dynamic defense frameworks this paper calls for. Further research could also explore perturbations in alternative acoustic domains, such as the cepstral domain or the raw waveform, to uncover new vulnerabilities and inform the design of more comprehensive, multi-domain defense strategies.
10. Conclusion
This study introduces the Spectral Masking and Interpolation Attack (SMIA), a novel black-box method that reveals critical vulnerabilities in modern voice security infrastructure. By strategically manipulating imperceptible frequency regions to naturalize synthetic audio, SMIA crafts adversarial samples that sound authentic yet consistently bypass defenses, achieving a 100% ASR against standalone countermeasures, 97.5% against voice authentication systems, and a strong 82% against combined VAS/CM pipelines. These findings conclusively demonstrate that current security postures, including layered defenses, are insufficient against adaptive threats. This research serves as an urgent call for a paradigm shift away from static models and toward dynamic, context-aware security frameworks capable of evolving with the threat landscape to protect the future of voice biometrics.
Ethical Considerations
This research introduces the SMIA to expose critical security gaps in modern VAS and CMs, with the primary objective of advancing defensive security research. The study demonstrates that current security postures are insufficient against adaptive adversarial attacks and highlights the urgent need for dynamic, context-aware defenses. By responsibly revealing these vulnerabilities, this work aims to assist service providers and researchers in developing next-generation security frameworks capable of detecting sophisticated deepfake threats.
All experiments were conducted under strict ethical guidelines to ensure data privacy and experimental safety. The evaluation relied exclusively on publicly available datasets, including ASVSpoof 2019 for benchmarking spoofing countermeasures (et al., 2019) and LibriSpeech for end-to-end system evaluation (Panayotov et al., 2015). No personal data was collected from real-world victims. Additionally, assessments involving commercial APIs (e.g., Microsoft Azure (Microsoft, )) were performed using these standardized datasets in controlled environments, ensuring no disruption to live services or unauthorized access to legitimate user accounts.
LLMs Usage Considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.